The Nonexistence of Certain Projective Planes of Order 15

David A. Drake Jean A. Larson

Abstract

If a projective plane Π of order 15 contains a line-extended (15,3)- Kirkman design Σ^{∗}, no collineation of Π of order 7 fixes (setwise) the point set of Σ^{∗}.